# FipTR: A Simple yet Effective Transformer Framework for Future Instance Prediction in Autonomous Driving

### v1

@@ Abstract
The future instance prediction from a Bird’s Eye View(BEV) perspective is a vital component in autonomous driving, which involves future instance segmentation and instance motion prediction. Existing methods usually rely on a redundant and complex pipeline which requires multiple auxiliary outputs and post-processing procedures. Moreover, estimated errors on each of the auxiliary predictions will lead to degradation of the prediction performance. In this paper, we propose a simple yet effective fully end-to-end framework named Future Instance Prediction Transformer(FipTR), which views the task as BEV instance segmentation and prediction for future frames. We propose to adopt instance queries representing specific traffic participants to directly estimate the corresponding future occupied masks, and thus get rid of complex post-processing procedures. Besides, we devise a flow-aware BEV predictor for future BEV feature prediction composed of a flow-aware deformable attention that takes backward flow guiding the offset sampling. A novel future instance matching strategy is also proposed to further improve the temporal coherence. Extensive experiments demonstrate the superiority of FipTR and its effectiveness under different temporal BEV encoders. The code will be released.
@@ 3.3 Flow-aware BEV Predictor
Unified future BEV queries. We predefined a group of grid-shaped learnable parameters $Q\in\mathrm{R}^{C\times H\times W}$ which represent the initial BEV feature for a future frame. We denote $Q_{t}^{p}$ as the BEV query located in a specific position $p=\{x,y\}$ at timestamp $t$. We share $Q$ across frames, i.e., $Q_{1}=Q_{2}=...=Q_{T_{out}}$. Following common practices [21], We add $Q_{t}$ with a learnable positional encoding, before feeding them into the following flow-aware deformable attention module to generate the updated BEV feature $F_{t}$.
Flow prediction module. Given the BEV feature map from the previous frame $F_{t-1}$, we use a flow prediction module to produce the backward flow map $\overleftarrow{f_{t}}$ which indicates the spatial correspondence relationship between the grids of $F_{t}$ and those of $F_{t_{1}}$ [25]. Concretely, $\overleftarrow{f_{t}}\in\mathbb{R}^{2\times H\times W}$ shares the same height $H$ and width $W$ with BEV feature map, while its two channels represent the offsets along the X- and Y-axis respectively. Take a specific grid with the coordinate $(x,y)$ in $F_{t}$ as an example, its corresponding grid in $F_{t-1}$ is the one with the coordinate $(x+\Delta x,y+\Delta y)$ and $(\Delta x,\Delta y)$ is the backward flow. In a similar way, we can establish the dense spatial correspondence relationship $\overleftarrow{f_{t}}$, which will be used to guide the offset sampling from $F_{t-1}$ to $F_{t}$.
@@
Before delving into the details of FADA, we first revisit the vanilla deformable attention(DA) mechanism, which takes $Q_{t}$ paired with its reference points $P$, and $F_{t-1}$ as inputs, and outputs $F_{t}$. We omit the subscript $t$ for clarity. Thus the calculation process can be formulated as
$\mathrm{DA}(Q_{p},p,F)=\\ \sum_{m=1}^{M}W_{m}\sum_{k=1}^{K}A_{mk}W^{\prime}_{m}F(p+\Delta p_{mk})$ (1)
where $p$ is a reference point and $Q_{p}$ is the corresponding query, respectively. $m$ indexes the attention head, and $M$ denotes the number of attention heads. $k$ indexes the sampled key and $K$ is the number of sampled keys for each head. $W_{m}\in\mathbb{R}^{C\times(C/M)}$ and $W_{m}^{\prime}\in\mathbb{R}^{(C/M)\times C}$ are learnable weights. $A_{mk}$ is the predicted attention weight. $\Delta p_{mk}$ is the sampling offsets for a specific reference point, and $F(p+\Delta p_{mk})$ refers to the input BEV feature at location $p+\Delta p_{mk}$.
@@ 3.4 Instance Masks Decoder
For future instance prediction, the segmentation head takes into consideration the instance queries and BEV feature map. To enrich the temporal cue of query and BEV feature at each timestamp, the updated instance query will be further transformed with a 3-layer Multi-layer Perceptron(MLP) and the BEV feature map will be transformed with a context subnetwork composed of two residual CNN blocks, and the parameters of all these modules are non-shared across frames, which are more adaptive and improve performance. The predicted masks are calculated by:
$\hat{m}_{t}^{i}=\mathrm{sigmoid}(Q_{I}^{t}\odot F_{t}^{\prime})$ (4)
where $\hat{m}_{t}^{i}\in\mathbb{R}^{1\times H\times W}$ is the mask logit, $Q_{I}^{t}$ and $F_{t}^{\prime}$ are the transformed instance queries and BEV feature map at timestamp $t\in\{0,1,...,T_{out}\}$, and $\odot$ is the Hadamard product, also known as element-wise product.
@@ 3.5 Future Instance Matching
Hungarian matching is an effective one-to-one matching strategy for end-to-end detection [4] or semantic segmentation [7] on a single frame. For tasks aiming at predicting a sequence of masks at consecutive frames, independently applying Hungarian matching to each frame cannot ensure the temporal coherence of instances.
@@ 4.1 Datasets and settings
We evaluate our approach on the NuScenes [3] dataset. Nuscenes contains 1000 scenes and each has 20 seconds annotated at 2HZ. For vision-centric methods, the provided sensory input includes six surrounding cameras with intrinsic and extrinsic calibration matrices and the ego-motions. We follow the training and evaluating settings used in previous works [12, 36, 1, 19] for fair comparison that the input is past one second states plus current state and the output is future two seconds predicted states. The BEV coordinate is based on the ego-vehicle system. The global prediction range is $100m\times 100m$ and the size of the generated BEV grid map is $200\times 200$, each of which corresponds to $0.5m\times 0.5m$.
For the model settings, we follow [36] which uses BEVDet4D [14] as the temporal BEV encoder and the backbone is SwinTransformer [24]. In addition, we also followed UniAD [13] which is based on BEVFormer [21] to verify the effectiveness and robustness of FipTR with different temporal BEV encoders. We set sampled key numbers $K=4$ and eight attention heads for four-layer flow-aware deformable attentions to predict future BEV feature sequences. We use six decoder layers of hidden dimension 256 and only the last layer is utilized to calculate the mask logits for computational saving.
@@
For inference, FipTR can fully predict the instance segmentation results for each frame simply using the result of the future masks decoder without any complex post-processing. Specifically, we multiply the object score and mask logits of a specific instance query as the final mask logits, $\hat{m}_{i}=c_{i}\cdot\hat{m}_{i}$ and set a threshold to filter which pixel belongs to the foreground $\hat{m}_{i}>\delta_{m}$.
@@ 4.4 Ablation Studies
Box and Mask Branch. Considering the heated discussion of the interrelationship between the box branch and mask branch [36, 31], we compare the detection and segmentation performance when removing the opposing branch (both the loss and matching cost) in Tab. 6. We can see that the auxiliary box branch plays an important role and directly predicting BEV masks in complex environments is difficult and the segmentation capability will be invalid. Moreover, We show the mean Average Precision (mAP) and mean Velocity Error (mAVE) which are 3D object detection metrics. Benefiting from the multi-task setting in matching cost and loss function, FipTR can conduct 3D object detection and future instance prediction simultaneously. The results indicate that the segmentation prediction is not negative to detection performance, and excitingly, the mAE metric has a significant improvement due to the motion modeling in FipTR.".

### v2

@@ Abstract
The future instance prediction from a Bird’s Eye View(BEV) perspective is a vital component in autonomous driving, which involves future instance segmentation and instance motion prediction. Existing methods usually rely on a redundant and complex pipeline which requires multiple auxiliary outputs and post-processing procedures. Moreover, estimated errors on each of the auxiliary predictions will lead to degradation of the prediction performance. In this paper, we propose a simple yet effective fully end-to-end framework named Future Instance Prediction Transformer(FipTR), which views the task as BEV instance segmentation and prediction for future frames. We propose to adopt instance queries representing specific traffic participants to directly estimate the corresponding future occupied masks, and thus get rid of complex post-processing procedures. Besides, we devise a flow-aware BEV predictor for future BEV feature prediction composed of a flow-aware deformable attention that takes backward flow guiding the offset sampling. A novel future instance matching strategy is also proposed to further improve the temporal coherence. Extensive experiments demonstrate the superiority of FipTR and its effectiveness under different temporal BEV encoders. The code is available at https://github.com/TabGuigui/FipTR.
@@ 3.3 Flow-aware BEV Predictor
Unified future BEV queries. We predefined a group of grid-shaped learnable parameters $Q\in\mathrm{R}^{C\times H\times W}$ which represent the initial BEV feature for a future frame. We denote $Q_{t,p}$ as the BEV query located in a specific position $p=\{x,y\}$ at timestamp $t$. We share $Q$ across frames, i.e., $Q_{1}=Q_{2}=...=Q_{T_{out}}$. Following common practices [21], we add $Q_{t}$ with a learnable positional encoding, before feeding them into the following flow-aware deformable attention module to generate the updated BEV feature $F_{t}$.
Flow prediction module. Given the BEV feature map from the previous frame $F_{t-1}$, we use a flow prediction module to produce the backward flow map $\overleftarrow{f_{t}}$ which indicates the spatial correspondence relationship between the grids of $F_{t}$ and those of $F_{t_{1}}$ [25]. Concretely, $\overleftarrow{f_{t}}\in\mathbb{R}^{2\times H\times W}$ shares the same height $H$ and width $W$ with BEV feature map, while its two channels represent the offsets along the X- and Y-axis respectively. Take a specific grid with the coordinate $(x,y)$ in $F_{t}$ as an example, its corresponding grid in $F_{t-1}$ is the one with the coordinate $(x+\Delta x,y+\Delta y)$ and $(\Delta x,\Delta y)$ is the backward flow. In a similar way, we can establish the dense spatial correspondence relationship $\overleftarrow{f_{t}}$, which will be used to guide the offset sampling from $F_{t-1}$ to $F_{t}$. The predicted flow map is supervised by the ground truth backward flow.
@@
Before delving into the details of FADA, we first revisit the vanilla deformable attention(DA) mechanism, which takes $Q_{t}$ paired with its reference point $p$ and $F_{t-1}$ as inputs, and outputs $F_{t}$. The calculation process can be formulated as:
$\mathrm{DA}(Q_{t,p},p,F_{t-1})=\\ \sum_{m=1}^{M}W_{m}\sum_{k=1}^{K}A_{mk}W^{\prime}_{m}F_{t-1}(p+\Delta p_{mk})$ (1)
where $p$ is a reference point and $Q_{t,p}$ is the corresponding query, respectively. $m$ indexes the attention head, and $M$ denotes the number of attention heads. $k$ indexes the sampled key and $K$ is the number of sampled keys for each head. $W_{m}\in\mathbb{R}^{C\times(C/M)}$ and $W_{m}^{\prime}\in\mathbb{R}^{(C/M)\times C}$ are learnable weights. $A_{mk}$ is the attention weight obtained via linear projection over the query and normalized by $\sum_{k=1}^{K}A_{mk}=1$. $\Delta p_{mk}$ is the sampling offsets for a specific reference point, and $F_{t-1}(p+\Delta p_{mk})$ refers to the input BEV feature at location $p+\Delta p_{mk}$.
@@ 3.4 Instance Masks Decoder
For future instance prediction, the segmentation head takes the instance queries and BEV feature as input. To enrich the temporal cue of query and BEV feature at each timestamp, the updated instance query will be transformed with a 3-layer Multi-layer Perceptron(MLP) and the BEV feature map will be transformed with a context subnetwork composed of two residual CNN blocks. The parameters of all these modules are non-shared across frames, which is more adaptive and improves performance. The predicted masks are calculated by:
$\hat{m}_{t}^{i}=\mathrm{sigmoid}(\mathrm{Q}_{i,t}\cdot F_{t}^{\prime})$ (4)
where $\hat{m}_{t}^{i}\in\mathbb{R}^{1\times H\times W}$ is the predicted mask logit, $\mathrm{Q}_{i,t}$ and $F_{t}^{\prime}$ are the transformed i-th instance query and BEV feature at timestamp $t\in\{0,1,...,T_{out}\}$.
@@ 3.5 Future Instance Matching
Hungarian matching is an effective one-to-one matching strategy for end-to-end detection [4, 40], semantic segmentation [7, 11] and tracking [36, 26]. For tasks aiming at predicting a sequence of masks at consecutive frames, independently applying Hungarian matching to each frame cannot ensure the temporal coherence.
@@ 4.1 Datasets and settings
We evaluate our approach on the NuScenes [3] dataset. Nuscenes contains 1000 scenes and each has 20 seconds annotated at 2HZ. For vision-centric methods, the provided sensory input includes six surrounding cameras with intrinsic and extrinsic calibration matrices and the ego-motions. We follow the training and evaluating settings used in previous works [12, 38, 1, 19] for fair comparison that the input is past one-second states plus current state and the output is future two seconds predicted states. The BEV coordinate is based on the ego-vehicle system. The global prediction range is $100m\times 100m$ and the size of the generated BEV grid map is $200\times 200$, each of which corresponds to $0.5m\times 0.5m$.
For the model settings, we follow [38] which uses BEVDet4D [14] as the temporal BEV encoder and the backbone is SwinTransformer [24]. In addition, we also followed UniAD [13] which is based on BEVFormer [21] to verify the effectiveness and robustness of FipTR with different temporal BEV encoders. The image size is set the same as the corresponding compared methods. We set sampled key numbers $K=4$ and eight attention heads for four-layer flow-aware deformable attentions to predict future BEV feature sequences. We use six decoder layers of hidden dimension 256 and only the last layer is utilized to calculate the mask logits for computational saving.
@@
For inference, FipTR predicts the instance segmentation results for each frame simply using the result of the future masks decoder without any complex post-processing. Specifically, we multiply the object score and mask logits of a specific instance query as the final mask logits, $\hat{m}_{i}=c_{i}\cdot\hat{m}_{i}$ and set a threshold to filter which grid belongs to the foreground $\hat{m}_{i}>\delta_{m}$. When a grid is occupied with multiple instances, we choose the instance with max logit.
@@ 4.4 Ablation Studies
Box and Mask Branch. Considering the heated discussion of the interrelationship between the box branch and mask branch [38, 32], we compare the detection and segmentation performance when removing the opposing branch (both the loss and cost) in Tab. 6. The auxiliary box branch plays an important role and directly predicting BEV masks in complex environments is difficult and the segmentation capability will be invalid. Moreover, we show the mean Average Precision (mAP) and mean Average Velocity Error (mAVE) which are 3D object detection metrics. Benefiting from the multi-task setting in matching cost and loss function, FipTR can conduct 3D object detection and future instance prediction simultaneously. The results indicate that the segmentation prediction is not negative to detection performance, and excitingly, the mAVE metric has a significant improvement due to the motion modeling in FipTR.